# Graph neural network outputs are almost surely asymptotically constant

### v1

@@ 6.2 Empirical results
In Figure 1, a single model initialization of the MeanGNN, GAT and GPS+RW architectures is experimented with over the different Erdős-Rényi graph distributions $\mathrm{ER}(n,p(n)=0.1)$, $\mathrm{ER}(n,p(n)=\frac{\log{n}}{n})$ and $\mathrm{ER}(n,p(n)=\frac{1}{50n})$. Each curve in the plots corresponds to a different class probability, depicting the average of 100 samples for each graph size along with the standard deviation shown in lower opacity.
@@
The convergence of class probabilities is apparent across all models and graph distributions, as illustrated in Figure 1, in accordance with our main theorems. The key differences between the plots are the convergence time, the standard deviation and the converged values.
@@
One striking feature of the plots in Figure 1 is that each given model converges to the same outputs on both the dense and the logarithmic growth distributions, while on the sparse distributions the eventual constant output varies. Our theorems do not predict this, but we can relate it to the distinct proof strategy employed in the sparse case, which uses convergence of the proportions of local isomorphism types. We can have many local isomorphism types, and the experiments show that what we converge to depends on the proportion of these. In the other cases, regardless of the distribution, the neighbourhood sizes are unbounded, so there is asymptotically almost surely one local graph type,
@@
As illustrated in Figure 1, it becomes apparent that attention-based models such as GAT and GPS+RW exhibit delayed convergence and greater standard deviation in comparison to MeanGNN. A possible explanation is that because some nodes are weighted more than others, the attention aggregation has a higher variance than regular mean aggregation. For instance, if half of the nodes have weights close to 00, then the attention aggregation effectively takes a mean over half of the available nodes. Eventually, however, the attention weights themselves converge, and thus convergence cannot be postponed indefinitely.
@@
Figure 2 depicts the outcomes of the GPS+RW architecture for various graph distributions. The analysis involves calculating the Euclidean distance between the class probabilities of the GPS+RW architecture and the mean class probabilities over the different graph samples. The standard deviation across the different graph samples is then derived for each of the 5 different model initializations and presented in Figure 2. The decrease of standard deviation across the different model initializations in Figure 2 indicates that all class probabilities converge across the different model initializations, empirically verifying the phenomenon for varying model initializations.
7 Discussion
@@
We have shown there is a wide convergence phenomenon for probabilistic classifiers expressed even in very advanced GNN architectures, and it applies to a great variety of random graph models. Rather than having separate proof techniques per GNN model, our paper introduces a broad language where such models can be situated, and provides techniques for showing convergence of term languages with respect to random graph models.
We believe that our approach extends to wider classes of distributions, such as variations of $\mathrm{SBM}$ with similar growth conditions as for ER, as well as distributions allowing dependence between features and graph structure. We leave the verification of these to future work. We are also investigating the case of non-weighted sum aggregation: here we believe that convergence laws can be obtained in the reals extended with positive and negative infinity.
We note that our term language can be extended with supremum and infimum operators. We can provide convergence laws for many Erdős-Rényi distributions using this extension, and this allows us to capture GNNs with max and min aggregation. And since the languages can already capture Boolean operators via Lipschitz functions, this extension subsumes classical first order logic. Thus our approach can subsume many convergence laws for predicate logic (Fagin, 1976). But we need more restrictive conditions on the edge probability function $p$ – and such restrictions are necessary, since it is easy to show that first-order logic does not obey a convergence law for Erdős-Rényi sparse random graphs, where sparsity is defined as in our Theorem 5.1 (Shelah & Spencer, 1988). Thus max aggregation gives an example not only of the interplay of our results with logical convergence laws, but of the boundary for convergence phenomena. We leave the description of the exact conditions for extending to max aggregation, as well as investigation of distributions for which asymptotic constancy fails, to future work.
Another important issue left for the future is an examination of convergence rates. In this work we only investigate them experimentally.
8 Social impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Concentration Inequalities
@@
Theorem A.1 (Markov’s Inequality).
@@
See Proposition 1.2.4, p. 8 of (Vershynin, 2018). ∎
Theorem A.2 (Chebyshev’s Inequality).
@@
See Corollary 1.2.5, p. 8 of (Vershynin, 2018). ∎
Theorem A.3 (Chernoff’s Inequality).
@@
See Theorem 2.3.1, p. 17 of (Vershynin, 2018) and Exercise 2.3.2, p. 18 of (Vershynin, 2018). ∎
Theorem A.4 (Hoeffding’s Inequality for bounded random variables).
@@
See Theorem 2.2.6, p. 16 of (Vershynin, 2018). ∎
Corollary A.5 (Hoeffding’s Inequality for Bernoulli random variables).
@@
Appendix B Proof for Erdős-Rényi distributions, the non-sparse cases
We new begin with the proof of convergence for the first three cases of the Erdős-Rényi distribution in Theorem 5.1.
@@ Appendix B Proof for Erdős-Rényi distributions, the non-sparse cases
B.1 Combinatorial and growth lemmas
Remark B.1.
@@ B.1 Combinatorial and growth lemmas
Note that in the first two cases, we only have claims about most tuples, while in the third case we make an assertion about most tuples, while also asserting an upper bound on all tuples. The upper bound on all tuples does not subsume the upper bound on most tuples, since the former states the existence of an $R_{3}$, while the latter gives a bound for an arbitrary $R_{2}$.
@@
    Now, any $d(u)$ is the sum of $n-1$ i.i.d. Bernoulli random variables with parameter $p(n)$. Hence by Hoeffding’s Inequality (Theorem A.4) and a union bound, there is $N_{2}$ such that for all $n\geq N_{2}$, with probability at least $1-\theta$, for every node $u$ we have:
@@
    Take $R_{1}\in(0,1)$ and $R_{2}>1$ and fix $\theta>0$. Again, any $d(u)$ is the sum of $n-1$ i.i.d. Bernoulli random variables with parameter $p(n)$. By Chernoff’s Inequality (Theorem A.3), for any node $u$ we have that:
@@
    Take $R_{1}\in(0,1)$ and $R_{2}>1$ and fix $\delta,\theta>0$. By Chernoff’s Inequality (Theorem A.3), for any node $u$ we have that:
@@
    By Markov’s Inequality (Theorem A.1) we have that:
@@
    To show the ‘moreover’ part, note that by Chernoff’s Inequality (Theorem A.3), for any $R_{3}>0$ and for any node $u$ we have that:
@@
Lemma B.3 (Regularity Lemma).
@@
By Lemma B.2 (1) there is $N$ such that for all $n\geq N$, with probability at least $1-\theta$, for all tuples $\bar{u}$ which satisfy $\phi(\bar{u})$ we have that:
@@
Applying Lemma B.4 with $\sigma=\delta+\delta^{2}$, $a=n(\tilde{p}-\epsilon)$ and $n=n(\tilde{p}+\epsilon)$, we get that:
@@
By Lemma B.2 (2) there is $N\in\mathbb{N}$ such that for all $n\geq N$, with probability at least $1-\theta$, for all tuples $\bar{u}$ which satisfy $\phi(\bar{u})$ we have that:
@@
By Lemma B.2 (3), there is $R_{3}>0$ and $N\in\mathbb{N}$ such that for all $n\geq N$, with probability at least $1-\theta$ when drawing graphs from $\mathrm{ER}(n,p(n))$, for at least $1-\delta$ of the tuples $\bar{u}$ which satisfy $\phi(\bar{u})$ we have that:
@@
This completes the proof of Lemma B.3 for the logarithmic growth case. ∎
@@
[…]
@@
We start with the Dense case. Let $p$ converge to $\tilde{p}>0$. Take $\epsilon,\delta,\theta>0$. By Hoeffding’s Inequality (Theorem A.4) and taking a union bound, there is $N$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that:
@@
As in the proof of Lemma B.3, by Chernoff’s Inequality (Theorem A.3) and taking a union bound, there are $0<R_{1}<1<R_{2}$ and $N$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that for all $v$:
@@
As in the proof of Lemma B.3, by Chernoff’s Inequality (Theorem A.3), there are $0<R_{1}<1<R_{2}$ and $N$ such that for all $n\geq N_{1}$ with probability at least $1-\theta$ we have that for at least $1-\delta$ proportion of nodes $v$:
@@
B.2 Proving the inductive invariant for the non-sparse cases, and proving the convergence theorem for these cases
@@
Now consider a term $\pi=\sum_{v}\rho(\bar{u},v)\star h(\eta(\bar{u},v))$. Define:
@@
$f_{\pi}(\bar{\bm{a}},t)\coloneqq\sum_{t^{\prime}\in\mathsf{Ext}(t)}\alpha_{u_{i}}(t,t^{\prime})\operatorname*{\mathbb{E}}_{{\bm{b}}\sim\mathbb{F}}\left[e_{\rho}(\bar{\bm{a}},{\bm{b}},t^{\prime})h(e_{\eta}(\bar{\bm{a}},{\bm{b}},t^{\prime}))\right]$
@@
$g_{\pi}(\bar{\bm{a}},t)\coloneqq\sum_{t^{\prime}\in\mathsf{Ext}(t)}\alpha_{u_{i}}(t,t^{\prime})\operatorname*{\mathbb{E}}_{{\bm{b}}\sim\mathbb{F}}\left[h(e_{\eta}(\bar{\bm{a}},{\bm{b}},t^{\prime}))\right]$
@@
  • Base case $\text{rw}(v)$. This follows by Lemma B.5.
@@
    By Lemma B.3 there is $N_{\gamma}>N_{1}$ such that for all $n\geq N_{\gamma}$, with probability at least $1-\theta$, out of all the tuples $\bar{u}$ such that $\psi(\bar{u})$, at least $1-\gamma$ satisfy that for at least $1-(\delta+\delta^{2})$ of the $v$’s satisfy equation ($\dagger_{\rho}$) and equation ($\dagger_{\eta}$).
    Now consider $t(\bar{u})\in\mathsf{GrTp}_{k}$ and take any $t^{\prime}(\bar{u},v)\in\mathsf{Ext}(t)$. Let $t^{\prime}$ specify that there are $r(t^{\prime})$ edges between $\bar{u}$ and $v$. For any $\bar{u}$ which satisfies $t$ the expected number of nodes $v$ which satisfy $t^{\prime}$ is:
    $p(n)^{r(t^{\prime})}(1-p(n))^{k-r(t^{\prime})}$
    Since $p(n)$ converges to $\tilde{p}$ and $\mathsf{Ext}(t)$ is finite, there is $N_{3}$ such that for all $n\geq N_{3}$ we have that for every pair $(t,t^{\prime})$:
    $|p(n)^{r(t^{\prime})}(1-p(n))^{k-r(t^{\prime})}-\alpha(t,t^{\prime})|<\frac{1}{2^{k+1}}\epsilon$ ($\ast$)
@@
    $f^{\circ}_{\pi}(v;\bar{u})\coloneqq e_{\rho}(\mathrm{H}(\bar{u},v),\mathrm{GrTp}(\bar{u},v))h(e_{\eta}(\mathrm{H}(\bar{u},v),\mathrm{GrTp}(\bar{u},v)))$
    Letting $t=\mathrm{GrTp}(\bar{u})$ have that:
    $\operatorname*{\mathbb{E}}_{v}[f^{\circ}_{\pi}(v;\bar{u})]=f_{\pi}(\mathrm{H}(\bar{u}),t)-\sum_{t^{\prime}\in\mathsf{Ext}(t)}p(n)^{r(t^{\prime})}(1-p(n))^{k-r(t^{\prime})}\operatorname*{\mathbb{E}}_{v}\left[e_{\rho}(\mathrm{H}(\bar{u},v),t^{\prime})h(e_{\eta}(\mathrm{H}(\bar{u},v),t^{\prime}))\right]$
    We note that:
    $f_{\pi}(\mathrm{H}(\bar{u}),t)$ $=\sum_{t^{\prime}\in\mathsf{Ext}(t)}\alpha_{u_{i}}(t,t^{\prime})\operatorname*{\mathbb{E}}_{{\bm{b}}\sim\mathbb{F}}\left[e_{\rho}(\mathrm{H}(\bar{u}),{\bm{b}},t^{\prime})h(e_{\eta}(\mathrm{H}(\bar{u}),{\bm{b}},t^{\prime}))\right]$
    $=\sum_{t^{\prime}\in\mathsf{Ext}(t)}\alpha_{u_{i}}(t,t^{\prime})\operatorname*{\mathbb{E}}_{v}\left[e_{\rho}(\mathrm{H}(\bar{u},v),t^{\prime})h(e_{\eta}(\mathrm{H}(\bar{u},v),t^{\prime}))\right]$
@@
    $\left\lVert f_{\pi}(\mathrm{H}(\bar{u}),\mathrm{GrTp}(\bar{u}))-\operatorname*{\mathbb{E}}_{v}[f^{\circ}_{\pi}(v;\bar{u})]\right\rVert<\frac{1}{2}\epsilon$ ($\square$)
    Furthermore note that the function:
    is bounded. Let $\Lambda$ be the diameter of its range. This also bounds $f^{\circ}_{\pi}$. Then by Hoeffding’s Inequality (Theorem A.4) for any $\bar{u}$, the probability that:
    $\left\lVert\frac{1}{n}\sum_{v}f^{\circ}_{\pi}(v;\bar{u})-\operatorname*{\mathbb{E}}_{v}[f^{\circ}_{\pi}(v;\bar{u})]\right\rVert\geq\frac{1}{2}\epsilon$
    is at most:
    $2d\exp\left(-\frac{\epsilon^{2}n}{2\Lambda^{2}}\right)$
    There is $N_{4}$ such that for all $n\geq N_{4}$ we have:
    $2d\exp\left(-\frac{\epsilon^{2}n}{2\Lambda^{2}}\right)<\theta\delta$
@@
    $\left\lVert\frac{1}{n}\sum_{v}f^{\circ}_{\pi}(v;\bar{u})-\operatorname*{\mathbb{E}}_{v}[f^{\circ}_{\pi}(v;\bar{u})]\right\rVert\geq\frac{1}{2}\epsilon$
@@
    $\left\lVert\frac{1}{n}\sum_{v}f^{\circ}_{\pi}(v;\bar{u})-\operatorname*{\mathbb{E}}_{v}[f^{\circ}_{\pi}(v;\bar{u})]\right\rVert<\frac{1}{2}\epsilon$
@@
    $\left\lVert\frac{1}{n}\sum_{v}f^{\circ}_{\pi}(v;\bar{u})-f_{\pi}(\mathrm{H}(\bar{u}),\mathrm{GrTp}(\bar{u}))\right\rVert<\epsilon$ ($\triangle_{f}$)
@@
    $g^{\circ}_{\pi}(v;\bar{u})\coloneqq h(e_{\eta}(\mathrm{H}(\bar{u},v),\mathrm{GrTp}(\bar{u},v)))$
@@
    $\left\lVert\frac{1}{n}\sum_{v}g^{\circ}_{\pi}(v;\bar{u})-g_{\pi}(\mathrm{H}(\bar{u}),\mathrm{GrTp}(\bar{u}))\right\rVert<\epsilon$ ($\triangle_{g}$)

### v2

@@ 6.1 Empirical results on random graph models
In Figure 2, a single model initialization of the MeanGNN, GAT and GPS+RW architectures is used with $\mathrm{ER}(n,p(n)=0.1)$, $\mathrm{ER}(n,p(n)=\frac{\log{n}}{n})$ and $\mathrm{ER}(n,p(n)=\frac{1}{50n})$. Each curve in the plots corresponds to a different class probability, depicting the average of 100 samples for each graph size along with the standard deviation shown in lower opacity.
Figure 3: Each plot depicts the standard deviation of Euclidean distances between class probabilities and their respective means across various samples of each graph size for GPS+RW.
@@
The convergence of class probabilities is apparent across all models and graph distributions, as illustrated in Figure 2, in accordance with our main theorems (Q1). The key differences between the plots are the convergence time, the standard deviation and the converged values.
@@
One striking feature of Figure 2 is that each given model converges to the same outputs on both the dense and the logarithmic growth distributions, while on the sparse distributions the eventual constant output varies (Q3). We can relate this to the distinct proof strategy employed in the sparse case, which uses convergence of the proportions of local isomorphism types. We can have many local isomorphism types, and the experiments show that what we converge to depends on the proportion of these. In the other cases, regardless of the distribution, the neighbourhood sizes are unbounded, so there is asymptotically almost surely one ‘local graph type’.
@@
We observe that attention-based models such as GAT and GPS+RW exhibit delayed convergence and greater standard deviation in comparison to MeanGNN (Q2). A possible explanation is that because some nodes are weighted more than others, the attention aggregation has a higher variance than regular mean aggregation. For instance, if half of the nodes have weights close to 00, then the attention aggregation effectively takes a mean over half of the available nodes. Eventually, however, the attention weights themselves converge, and thus convergence cannot be postponed indefinitely.
@@
Figure 3 depicts the outcomes of the GPS+RW architecture for various graph distributions. The analysis involves calculating the Euclidean distance between the class probabilities of the GPS+RW architecture and the mean class probabilities over the different graph samples. The standard deviation across the different graph samples is then derived for each of the 5 different model initializations and presented in Figure 3. The decrease of standard deviation across the different model initializations in Figure 3 indicates that all class probabilities converge across the different model initializations, empirically verifying the phenomenon for varying model initializations (Q1 and Q3).
6.2 Empirical results on large real-world graphs
Figure 4: Standard deviation of distances between class probabilities and their means across TIGER-Alaska graph sizes for MeanGNN.
Towards Q4, we investigated a large real-world dataset. Many commonly studied graph datasets (e.g. ZINC [13], QM9 [6]) do not exhibit sufficient graph size variance and provide no obvious means to add scaling. We used the TIGER-Alaska dataset [16] of geographic faces. The original dataset has 93366 nodes, while Dimitrov et al. [10] extracted smaller datasets with graphs having 1K, 5K, 10K, 25K and 90K nodes. We chose the modified dataset as it is split by graph size, and consists of graphs differing from our random models (in particular all graphs are planar). Figure 4 shows the results of applying the same five MeanGNNs to graphs of increasing sizes. Strikingly, we again observe a convergence phenomenon, but at a slower pace.
7 Discussion and limitations
@@
We have shown there is a wide convergence phenomenon for probabilistic classifiers expressed even in very advanced GNN architectures, and it applies to a great variety of random graph models. Rather than having separate proof techniques per GNN model, our paper introduces a broad language where such models can be situated, and provides techniques at the level of term languages.
We emphasize that a.a.s. convergence is not universal, and it does not hold even for ER with arbitrary $p(n)$ going to 00. It is easy to see that if we alternate between the sparse case and the log or root case, we will get divergence. A.a.s. convergence also depends strongly on the details of the term language: if we add sum aggregation, the term values clearly diverge, and similarly if we allow non-smooth functions – e.g. testing whether a term is strictly above a constant. A.a.s. convergence does not hold if max aggregation is added, although we believe that weaker convergence laws can be established for this aggregator.
Appendix A Graph Convolutional Networks
In the body of the paper, we mentioned that our results can be extended to graph convolutional networks (GCNs) [29]. In this section we show how to extend our term language to capture them. Later in the appendix when providing proof details for our convergence laws, we will utilize the extended language, thus showing the applicability to GCNs.
GCNs are instances of the message passing neural network in which the update function is defined as follows.
${\bm{h}}_{v}^{\left(\ell+1\right)}=\sum_{u\in\operatorname{\mathcal{N}}(v)}\frac{1}{\sqrt{|\operatorname{\mathcal{N}}(v)||\operatorname{\mathcal{N}}(u)|}}{\bm{W}}^{(\ell)}{\bm{h}}_{u}^{(\ell)}$
To allow our term language to capture GCNs, we add the new aggregator:
${\textsc{GCN}}_{y\in\operatorname{\mathcal{N}}(x)}\tau(y)$
[…]
$\left\llbracket{\textsc{GCN}}_{y\in\operatorname{\mathcal{N}}(x_{i})}\tau(\bar{u})\right\rrbracket_{G}=\sum_{v\in\operatorname{\mathcal{N}}(u_{i})}\frac{1}{\sqrt{|\operatorname{\mathcal{N}}(u_{i})||\operatorname{\mathcal{N}}(v)|}}\llbracket\tau(\bar{u},v)\rrbracket_{G}$
With this operator it is possible to capture GCNs in the same way as MeanGNNs are captured (Section 4). Moreover, the extended term language permits arbitrary combinations of the GCN aggregator with other language features.
Appendix B Concentration Inequalities
@@
Theorem B.1 (Markov’s Inequality).
@@
See Proposition 1.2.4, p. 8 of [53]. ∎
Theorem B.2 (Chebyshev’s Inequality).
@@
See Corollary 1.2.5, p. 8 of [53]. ∎
Theorem B.3 (Chernoff’s Inequality).
@@
See Theorem 2.3.1, p. 17 of [53] and Exercise 2.3.2, p. 18 of [53]. ∎
Theorem B.4 (Hoeffding’s Inequality for bounded random variables).
@@
See Theorem 2.2.6, p. 16 of [53]. ∎
Corollary B.5 (Hoeffding’s Inequality for Bernoulli random variables).
@@
Appendix C Proof for Erdős-Rényi distributions, the non-sparse cases
We new begin with the proof of convergence for the first three cases of the Erdős-Rényi distribution in Theorem 5.1, which we restate here for convenience:
Theorem C.1.
Consider $(\mu_{n})_{n\in\mathbb{N}}$ sampling a graph $G$ from the Erdős-Rényi distribution $\mathrm{ER}(n,p(n))$ and node features independently from i.i.d. bounded distributions on $d$ features, where $p$ satisfies any of the following properties.
  • Density. $p$ converges to $\tilde{p}>0$.
  • Root growth. For some $K>0$ and $10<\beta<$ 1 we have: $p(n)=Kn^{-\beta}$.
  • Logarithmic growth. For some $K>0$ we have: $p(n)=K\frac{\log(n)}{n}$.
[…]
@@ Appendix C Proof for Erdős-Rényi distributions, the non-sparse cases
C.1 Combinatorial and growth lemmas
Remark C.2.
@@ C.1 Combinatorial and growth lemmas
Note that in the first two cases, we only have claims about all tuples, while in the third case we make an assertion about most tuples, while also asserting an upper bound on all tuples. The upper bound on all tuples does not subsume the upper bound on most tuples, since the former states the existence of an $R_{3}$, while the latter gives a bound for an arbitrary $R_{2}$.
@@
    Now, any $d(u)$ is the sum of $n-1$ i.i.d. Bernoulli random variables with parameter $p(n)$. Hence by Hoeffding’s Inequality (Theorem B.4) and a union bound, there is $N_{2}$ such that for all $n\geq N_{2}$, with probability at least $1-\theta$, for every node $u$ we have:
@@
    Take $R_{1}\in(0,1)$ and $R_{2}>1$ and fix $\theta>0$. Again, any $d(u)$ is the sum of $n-1$ i.i.d. Bernoulli random variables with parameter $p(n)$. By Chernoff’s Inequality (Theorem B.3), for any node $u$ we have that:
@@
    Take $R_{1}\in(0,1)$ and $R_{2}>1$ and fix $\delta,\theta>0$. By Chernoff’s Inequality (Theorem B.3), for any node $u$ we have that:
@@
    By Markov’s Inequality (Theorem B.1) we have that:
@@
    To show the ‘moreover’ part, note that by Chernoff’s Inequality (Theorem B.3), for any $R_{3}>0$ and for any node $u$ we have that:
@@
Lemma C.4 (Regularity Lemma).
@@
By Lemma C.3 (1) there is $N$ such that for all $n\geq N$, with probability at least $1-\theta$, for all tuples $\bar{u}$ which satisfy $\phi(\bar{u})$ we have that:
@@
Applying Lemma C.5 with $\sigma=\delta+\delta^{2}$, $a=n(\tilde{p}-\epsilon)$ and $n=n(\tilde{p}+\epsilon)$, we get that:
@@
By Lemma C.3 (2) there is $N\in\mathbb{N}$ such that for all $n\geq N$, with probability at least $1-\theta$, for all tuples $\bar{u}$ which satisfy $\phi(\bar{u})$ we have that:
@@
By Lemma C.3 (3), there is $R_{3}>0$ and $N\in\mathbb{N}$ such that for all $n\geq N$, with probability at least $1-\theta$ when drawing graphs from $\mathrm{ER}(n,p(n))$, for at least $1-\delta$ of the tuples $\bar{u}$ which satisfy $\phi(\bar{u})$ we have that:
@@
This completes the proof of Lemma C.4 for the logarithmic growth case. ∎
@@
$\left\lVert\text{rw}_{k}(v)\right\rVert<\epsilon$
@@
We start with the Dense case. Let $p$ converge to $\tilde{p}>0$. Take $\epsilon,\delta,\theta>0$. By Hoeffding’s Inequality (Theorem B.4) and taking a union bound, there is $N$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that:
@@
As in the proof of Lemma C.4, by Chernoff’s Inequality (Theorem B.3) and taking a union bound, there are $0<R_{1}<1<R_{2}$ and $N$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that for all $v$:
@@
As in the proof of Lemma C.4, by Chernoff’s Inequality (Theorem B.3), there are $0<R_{1}<1<R_{2}$ and $N$ such that for all $n\geq N_{1}$ with probability at least $1-\theta$ we have that for at least $1-\delta$ proportion of nodes $v$:
@@
C.2 Proving the inductive invariant for the non-sparse cases, and proving the convergence theorem for these cases
With the growth lemmas for the non-sparse cases in hand, we return to presenting the main proof of Theorem C.1, the convergence theorem for non-sparse ER distributions.
@@
Now consider a term $\pi=\sum_{v}\rho(\bar{u},v)\star h(\eta(\bar{u},v))$.
Recall that the semantics are:
$\frac{\sum\limits_{v\in G}\llbracket\rho(\bar{u},v)\rrbracket_{G}h(\llbracket\eta(\bar{u},v)\rrbracket_{G})}{\sum\limits_{v\in G}h(\llbracket\eta(\bar{u},v)\rrbracket_{G})}$
Define:
@@
$f_{\pi}(\bar{\bm{a}},t)\coloneqq\sum_{t^{\prime}\in\mathsf{Ext}_{u_{i}}(t)}\alpha_{u_{i}}(t,t^{\prime})\operatorname*{\mathbb{E}}_{{\bm{b}}\sim\mathbb{F}}\left[e_{\rho}(\bar{\bm{a}},{\bm{b}},t^{\prime})h(e_{\eta}(\bar{\bm{a}},{\bm{b}},t^{\prime}))\right]$
@@
$g_{\pi}(\bar{\bm{a}},t)\coloneqq\sum_{t^{\prime}\in\mathsf{Ext}_{u_{i}}(t)}\alpha_{u_{i}}(t,t^{\prime})\operatorname*{\mathbb{E}}_{{\bm{b}}\sim\mathbb{F}}\left[h(e_{\eta}(\bar{\bm{a}},{\bm{b}},t^{\prime}))\right]$
@@
  • Base case $\text{rw}(v)$. This follows by Lemma C.6.
@@
    By Lemma C.4 there is $N_{\gamma}>N_{1}$ such that for all $n\geq N_{\gamma}$, with probability at least $1-\theta$, out of all the tuples $\bar{u}$ such that $\psi(\bar{u})$, at least $1-\gamma$ satisfy that for at least $1-(\delta+\delta^{2})$ of the $v$’s satisfy equation ($\dagger_{\rho}$) and equation ($\dagger_{\eta}$).
    Now consider $t(\bar{u})\in\mathsf{GrTp}_{k}$ and take any $t^{\prime}(\bar{u},v)\in\mathsf{Ext}(t)$. For any tuple $\bar{u}$ of nodes such that $\mathrm{GrTp}(\bar{u})=t$, define:
    $\llbracket t^{\prime}(\bar{u})\rrbracket\coloneqq\{v\mid t^{\prime}(\bar{u},v)\}$
    Since $p(n)$ converges to $\tilde{p}$ and $\mathsf{Ext}(t)$ is finite, there is $N_{3}$ such that for all $n\geq N_{3}$ with probability at least $1-\theta$ we have that for every pair $(t,t^{\prime})$ and tuple $\bar{u}$ of nodes such that $t(\bar{u})$:
    $\left|\frac{|\llbracket t^{\prime}(\bar{u})\rrbracket|}{n}-\alpha(t,t^{\prime})\right|<\frac{1}{2^{k+1}}\epsilon$ ($\ast$)
@@
    $f^{\circ}_{\pi}(\bar{u},v)\coloneqq e_{\rho}(H_{G}(\bar{u},v),\mathrm{GrTp}(\bar{u},v))h(e_{\eta}(H_{G}(\bar{u},v),\mathrm{GrTp}(\bar{u},v)))$
    Note that the function:
    is bounded. Let $\Lambda$ be the diameter of its range. This also bounds $f^{\circ}_{\pi}$.
    We will now apply Hoeffding’s Inequality to $\sum_{v\mid t^{\prime}(\bar{u},v)}f^{\circ}_{\pi}(\bar{u},v)$. Note that the number of summands is $|\llbracket t^{\prime}(\bar{u})\rrbracket|$ and the summands are bounded by $\Lambda$. Furthermore, the summands are independent and have expected value:
    $\operatorname*{\mathbb{E}}_{{\bm{b}}\sim\mathbb{F}}\left[e_{\rho}(H_{G}(\bar{u},{\bm{b}}),t^{\prime})h(e_{\eta}(H_{G}(\bar{u},{\bm{b}}),t^{\prime}))\right]$
    Hence, by Hoeffding’s Inequality (Theorem B.4) for any $\bar{u}$ the probability that:
    $\left\lVert\frac{1}{n}\sum_{v\mid t^{\prime}(\bar{u},v)}f^{\circ}_{\pi}(\bar{u},v)-\frac{|\llbracket t^{\prime}(\bar{u})\rrbracket|}{n}\operatorname*{\mathbb{E}}_{{\bm{b}}\sim\mathbb{F}}\left[e_{\rho}(H_{G}(\bar{u},{\bm{b}}),t^{\prime})h(e_{\eta}(H_{G}(\bar{u},{\bm{b}}),t^{\prime}))\right]\right\rVert\geq\frac{1}{2^{k+1}}\epsilon$
    is at most:
    $2d\exp\left(-\frac{\epsilon^{2}|\llbracket t^{\prime}(\bar{u})\rrbracket|}{2\Lambda^{2}}\right)$
@@
    $2d\exp\left(-\frac{\epsilon^{2}|\llbracket t^{\prime}(\bar{u})\rrbracket|}{2\Lambda^{2}}\right)<\theta\delta$
@@
    $\left\lVert\frac{1}{n}\sum_{v\mid t^{\prime}(\bar{u},v)}f^{\circ}_{\pi}(\bar{u},v)-\frac{|\llbracket t^{\prime}(\bar{u})\rrbracket|}{n}\operatorname*{\mathbb{E}}_{{\bm{b}}\sim\mathbb{F}}\left[e_{\rho}(H_{G}(\bar{u},{\bm{b}}),t^{\prime})h(e_{\eta}(H_{G}(\bar{u},{\bm{b}}),t^{\prime}))\right]\right\rVert\geq\frac{1}{2^{k+1}}\epsilon$
@@
    $\left\lVert\frac{1}{n}\sum_{v\mid t^{\prime}(\bar{u},v)}f^{\circ}_{\pi}(\bar{u},v)-\frac{|\llbracket t^{\prime}(\bar{u})\rrbracket|}{n}\operatorname*{\mathbb{E}}_{{\bm{b}}\sim\mathbb{F}}\left[e_{\rho}(H_{G}(\bar{u},{\bm{b}}),t^{\prime})h(e_{\eta}(H_{G}(\bar{u},{\bm{b}}),t^{\prime}))\right]\right\rVert<\frac{1}{2^{k+1}}\epsilon$
@@
    $\left\lVert\frac{1}{n}\sum_{v}f^{\circ}_{\pi}(\bar{u},v)-f_{\pi}(H_{G}(\bar{u}),\mathrm{GrTp}(\bar{u}))\right\rVert<\epsilon$ ($\triangle_{f}$)
@@
    $g^{\circ}_{\pi}(\bar{u},v)\coloneqq h(e_{\eta}(H_{G}(\bar{u},v),\mathrm{GrTp}(\bar{u},v)))$
@@
    $\left\lVert\frac{1}{n}\sum_{v}g^{\circ}_{\pi}(\bar{u},v)-g_{\pi}(H_{G}(\bar{u}),\mathrm{GrTp}(\bar{u}))\right\rVert<\epsilon$ ($\triangle_{g}$)